HUBBLE SPACE TELESCOPE DAILY REPORT #1745 PERIOD COVERED: 0000Z (UTC) 10/23/96 - 0000Z (UTC) 10/24/96 Daily Status Report as of 298/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of FGS 6919 (3 Points of Light Calibration) The FGSs were used to make a Three Points of Light calibration by = making observations of standard star UPGREN69. This calibration will be = done at three different points in FGS#3 using three different filters = (pupil, clear and neutral density). This will be used to obtain a = reference transfer function which will be used to deduce the presence of = secondary stars in multiple star systems. The observations were executed = as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.3 Completed FOS/RD 6203 (FOS Cycle 5: Photometric Calibration) The Faint Object Spectrograph (Red Side) was used to make = observations of standard star BD+26D4211 in order to perform an absolute = sensitivity photometric calibration of the instrument. Following a = successful guide star acquisition at 297/061550Z, there was an FOS status = buffer message indicating that the on-board peakup real-time command = sequence timed out while waiting for the HST to become stable after the = slew to the dwell point. As a result the take data flag did not come up. = HSTAR 5152 was written. Otherwise, the observations executed as = scheduled, and no further anomalies were noted. 1.4 Completed WF/PC-2 7203 (Cosmology with the Medium Deep Survey - = Short Orbit Cases) The WF/PC-2 was used to take short (about 20-minute) exposures, to = be taken in parallel with the prime science observations, using the I = filter. These exposures are part of the medium-deep survey, which = emphasizes exposures that will be useful in obtaining cosmological = parameters. The observations were completed as planned, and no problems = were reported. 1.5 Completed GHRS/1 6454 (UV Spectroscopy of a DAO Dwarf Showing = Signatures of an Extremely Hot Wind) The High Resolution Spectrograph (Side One) was used to make = ultraviolet spectrographic observations of the DAO peculiar white dwarf = star HS2115+114 which is one of the small group of stars in the = Hamburg-Schmidt Survey. The spectra of this class of stars exhibit = shallow metal absorbing lines from ultra-high excited ionization stages. = The observations were executed as scheduled, and no anomalies were noted. = 1.6 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.7 Completed FOS/BL 6682 (Geometry and Generalizability of the = Reflected Light Model for Seyfert Galaxies) The Faint Object Spectrograph (Blue Side) was used to make = spectropolarimetry observations of the Seyfert (class two) galaxy MARK78 = in order to study its geometry and the generalizablity of its reflected = light model. The goal is to determine if the polarization components of = all Seyfert class two galaxies have Seyfert class one characteristics. = The observations were completed as scheduled, and no problems were = reported. 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: There were no reported operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH